OpenStudy (mathhelp346):

Write an equation for the line that is perpendicular to the line x-3y=9 at the x-intercept.

OpenStudy (anonymous):

What is the slope of a line perpendicular to the line x-3y=9 ?

OpenStudy (mathhelp346):

3

OpenStudy (anonymous):

right, and if you know the line goes through the x-intercept of x-3y=9, then you can find the y-coord of that intercept. Then you have a point and a slope of the perpendicular line, and you should be able to write the equation.

OpenStudy (mathhelp346):

y=3x-3?

OpenStudy (mathhelp346):

is that correct?

OpenStudy (anonymous):

I'm working it, just a sec

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

That's what I got, but I did it fast... might have made an error, so double check

OpenStudy (mathhelp346):

how did u get y=-3x+27?

OpenStudy (anonymous):

make original equation into y=mx+b x-3y=9 3y=x-9 y=(1/3)x - 3 The slope is (1/3), and the y intercept is (0,-3). I also found the x intercept at (9,0) Then the perp. line has a slope that is the negative inverse, or - 1/(1/3) = -3 I then used the x intercept value as a point and the slope of -3 to get the equation of the perp line

OpenStudy (anonymous):

The reason I used the x intercept for the point is that the problem said the perp line goes through that point

OpenStudy (mathhelp346):

so the equation would be y=-3x-3?

OpenStudy (anonymous):

The x intercept point from the original line is (9,0). That point isn't on the line equation you just wrote.

OpenStudy (anonymous):

you got the slope right, but the intercept wrong.

OpenStudy (anonymous):

I thought to myself, "hmmm... perp line must be y = -3x + b, and I knew that (9,0) was on the line, so I put in x=9 and y = 0 and found that b must be = 27."

OpenStudy (mathhelp346):

ohh i see

OpenStudy (anonymous):

I gotta run... sorry to leave abruptly, but hope this is helping...

OpenStudy (mathhelp346):

thanks

OpenStudy (anonymous):

glad to help :) good luck!